please help me solve this equation: (b+1/3=2
Answer by ptaylor(2198)


b+1/3=2 First, subtract 1/3 from both sides
b+1/3-1/3=2-1/3 Now we know that 2=6/3, right? And we also know that +1/3-1/3=0 So we have
b+0=6/3-1/3 and this equals 5/3
So b=5/3
Ck
5/3+1/3=2
6/3=2
2=2

Hope this helps----ptaylor